A father is thrice as old as his son. Five years ago, he was four times as old. How old is the father?
Answer by macston(5194)

f=father's age; s=son's age
f=3s
f-5=4(s-5) Substitute for f.
3s-5=4s-20 Add 20 to each side.
3s+15=4s Subtract 3s from each side.
15=s The son is 15 years old.
f=3s=3(15)=45 ANSWER: The father is 45 years old.
CHECK:
f-5=4(s-5)
45-5=4(15-5)
40 yrs=4(10 yrs)
40 yrs=40 yrs